OpenStudy (anonymous):

the sum of a number and its reciprocal is 13/6. find the numbers.

OpenStudy (kropot72):

Let the number = x \[x+\frac{1}{x}=\frac{13}{6}\] \[\frac{x ^{2}+1}{x}=\frac{13}{6}\] \[x ^{2}+1=\frac{13}{6}x\] \[x ^{2}-\frac{13}{6}+1=0\ ................(1)\] Now solve quadratic (1) using the formula that will solve any quadratic.